Please help! Thank you in advance!
Two ships A and B start from the same port at 8am. Ship A heads north 62 degrees 15 minutes east with a speed of 12 km/hr. Ship B takes the course of north 48 degrees 20 minutes west with the speed of 16 km/hr. How far apart are they at 10:45am? What is the bearing of B from A?

Answer by stanbon(75887)

Two ships A and B start from the same port at 8am. Ship A heads north 62 degrees 15 minutes east with a speed of 12 km/hr. Ship B takes the course of north 48 degrees 20 minutes west with the speed of 16 km/hr. How far apart are they at 10:45am? What is the bearing of B from A?
----
Draw the picture.
Distance for A = side A = rate*time = 12*2.75 = 30 km
Distance for B = side B = rate*time = 16*2.75 = 40 km
------
Angle made by the two sides = C = 62.25+48.33 = 110.58 degrees
------
Solve for side "c" using the Law of Cosines.
c^2 = a^2 + b^2-2ab*cos(C)
-----
c = sqrt(3343.63) = 57.82 km
==================================
Bearing of B from A = 110.58 degrees
----------------------------------
Cheers,
Stan H
----